# Ultrafast Solvent Dynamics Drives the Formation of the Hydrated Electron in Photoexcited Water

### v1

@@ 4.2 Excited-State Molecular Dynamics
The excited state molecular dynamics (ESMD) simulations were initiated from the first electronic excited state ($S_{1}$), computed using the Restricted Open-shell Kohn–Sham (ROKS) method 44, 45. ROKS has proven to be a reliable approach for describing dissociative and charge-transfer states in closed-shell systems 46, 47, 48, processes that play a central role in the photochemistry of water.
@@
To propagate the system on the $S_{1}$ potential energy surface, we employed the Atomic Simulation Environment (ASE)94, a python module which is interfaced with CP2K software95. At each timestep, we calculated the energy on the electronic ground state $S_{0}$ using Density Functional Theory (DFT) to account for possible non-radiative relaxation pathways49, 47 (see below) which in our current setup is determined by examining the energy gap. The simulations were performed in the ensemble NVE using a timestep of 0.5 fs and velocity verlet algorithm to perform the nuclei propagation of the system.
@@
Both excited and ground state calculations were performed using the hybrid PBEh(40)-rVV10 functional96 in combination with the DZVP-MOLOPT-SR basis set97. This level of electronic structure theory has been successfully applied to the study of liquid water98, 99, 100, demonstrating its accuracy in reproducing ionization potential, redox potentials, electronic levels of excess electron. Carter et. al49. also reported excellent agreement with experiments for aqueous iodine solutions using the same functional within the ROKS framework. More recently the functional PBEh(40) was identified as one of the most reliable choices for describing the hydrated electron in its electronic ground state68. In our current work, the favorable agreement with reproducing the experimental UV-absorption and other quantities sensitive to the excited states of liquid water gives us confidence in the quality of our results.
@@
Each ESMD simulation was propagated until the system reached a crossing point, defined as the step where the energy gap between the first excited state ($S_{1}$) and the ground state ($S_{0}$) dropped below 0.2 eV. This criterion provides a robust and efficient way to identify non-radiative decay events and allows us to characterize the distribution of deactivation pathways, offering a lower bound on the excited-state lifetimes61, 62. A known limitation of this approach is that it does not explicitly include non-adiabatic coupling vectors (NACVs), which are key quantities for describing electronic transitions during non-radiative decay. However, a previous study on the radiolysis of water cluster in vacuum showed that NACVs increase significantly as the energy gap decreases101. This strong correlation implies that small energy gaps are reliable indicators of high non-adiabatic transition probabilities. Therefore, by using the energy gap as a decay criterion, we effectively capture the relevant non-radiative dynamics, justifying the omission of explicitly determining NACVs in our simulations.
@@
The simulations that did not exhibit a decay were continued until a maximum time of 1 ps was reached. Spin density cube files were calculated and outputted every 5 fs during all ESMD trajectories and at the step of crossing, for post-processing analysis (see section below). All the electronic calculations in the present work were performed using CP2K version 2024.295.
@@ 5.1 Coordination Numbers
Here, $i$ is the atom index, while $\kappa=10$, $\textbf{r}_{c}=1.38$ Å are parameters taken from previous works used to study the dissociation of water in the ground-state to form hydronium and hydroxide ions102, 103, 104, 105. In this manner, for an isolated hydrogen atom not attached to any water molecule, the $CN_{H}$ is zero while for a hydrogen covalently bonded to a water molecule $CN_{H}$ is 1. In the case of $CN_{O}$, it takes on a value of 3, 2 and 1, for a hydronium (ion or radical), neutral molecular water and hydroxyl species respectively.
@@ 5.2 Hydrogen Bond (HB) Network Classification
HB were identified using a geometrical criteria introduced by Kumar et. al.106: O-O distance of $\leq 3.5$ Å and H-O—O angle $\leq 30\ deg$. A water molecule with four HB is classified as part of a normal HB network, otherwise, it is labeled as a defect (see Figure 1C for categories). Sensitivity analysis to the choice of these geometric parameters is addressed in Section S3 of the Supplementary Information showing that while the hydrogen-bond criteria affect the exact populations of defects, the trends and physical insights inferred are robust.
@@ 5.3 Spin-Density Analysis for Excitation Localization
where $r_{0}^{i}$ is the position of the ithi-th oxygen atom and $R=1$ Å. To estimate the degree of localization, we computed the Inverse Participation Ratio (IPR)51, 50:
@@ 5.4 HAT versus PCET Mechanism
A direct calculation of the electron center from the spin density cube files yields a position located between the $\mathrm{HO}^{\bullet}$ and the $\mathrm{H}^{\bullet}$ (Equation 4), or between the $\mathrm{HO}^{\bullet}$ and the excess electron (Equation 5). Figure S5 in the Supplementary Information illustrates this point. This result lacks a clear physical interpretation, as it does not correspond to the actual location of any species and essentially implies that the spin density is located close to the geometric center of the species involved. This problem does not arise in earlier theoretical studies27, 24, 43, 70, as those simulations typically start with a pre-inserted electron in bulk water and the formation of an additional radical species are not observed.
@@
To overcome this challenge, we first identify the $\mathrm{HO}^{\bullet}$ species by locating the water molecule with the longest OH bond in each trajectory (see Section S6 of the Supplementary Information). Once identified, we remove the contribution of the $\mathrm{HO}^{\bullet}$ to the spin density by excluding a spherical region of radius 1.5 Å centered on its oxygen atom (see Section S7 in the Supplementary Information). This step yields a modified cube file that provides a more meaningful representation of the spin density to localize the hydrated excess electron. We then compute the electron center and gyration radius from this file using the protocol described in previous studies43 (see Section S8 in the Supplementary Information). Since the cube files were saved every 5 fs and at the crossing geometry, the electron center and gyration radius for the intermediate frames were obtained using linear interpolation between consecutive cube files.
@@
Once all the species were correctly identified, we classified each trajectory according to its reaction mechanism. For this, we analyzed the final frame of each simulation (corresponding to the geometry at the point of non-radiative crossing) and we computed the distance between the electron center and the nearest hydrogen atom. If the distance was less than $0.6$ Å, the trajectory was classified as hydrogen atom transfer (HAT), otherwise it was assigned to the proton coupled electron transfer (PCET). The $0.6$ Å threshold was chosen based on the radial distribution function (RDF) between the electron and all hydrogen atoms (see Figure S9 in the Supplementary Information). Specifically, it represents the midpoint between two cases: a bound hydrogen atom (distance electron and nearest H $\sim$0 Å) and the first maximum associated with hydrogen atoms in the solvation shell of the electron (distance between electron and nearest H $\sim$1.2 Å)43, 49. A short electron–hydrogen distance indicates that both species occupy the same spatial region, consistent with the formation of a neutral hydrogen atom. In contrast, larger distances suggest that the electron remains solvated while the proton is bound to nearby water molecules, consistent with a PCET process.
@@
Figure S1: Initial photo-absorption
@@
This partitioning of the spin density is a key variable used in our approach to identify the water molecules that are involved in the photoexcitation in an agnostic manner. However, this approach rely on a parameter, which is the cutoff radius. Figure S2 shows the IPR histograms (same to the one presented in Figure 1B in the main text) using different parameters. We can clearly see that numerical values are different however they all follow the same physical trends: most of the excitation are localized on a single water molecules and a small contribution of the delocalized excitation can be observed.
@@
Figure S4: Decay times
@@
In this section, we address one of the key challenges discussed in the manuscript, the proper identification of the released electron. Figure S5 presents the spin densities for two representative frames, each corresponding to one of the two mechanisms observed in the photochemistry of water.
Figure S5: Identification of released electron. Original cubes files are shown in the left panels (as obtained from the CP2K software using ROKS method). The Modified cube files obtained as an output of our procedure are shown in the right panels. The center of the spin density for each kind of cube files are represented as a blue ball. We show the cube files for a selected frame exhibiting both PCET (upper panels) and HAT (lower panels) mechanisms.
As shown in the left panels of Figure S5, calculating the electron center using the original cube files results in a position (indicated by a light-blue sphere) located approximately in the middle between the $\mathrm{HO}^{\bullet}$ and the hydrated electron (top-left panel) or the $\mathrm{H}^{\bullet}$ (bottom-left panel). To address this issue, our protocol, described in the manuscript, generates a modified cube file. This improved representation, shown in the right panels of Figure S5, allows for a more reliable calculation of properties such as the electron center and the gyration radius.
@@
In all trajectories, we observed that the dissociating species originates from the water molecule involved in the initial photo-absorption event. However, in cases where the excitation is delocalized (see Figure 1B in the main manuscript), it is not straightforward to determine in advance which specific molecule will dissociate. To automate this identification, we analyzed all trajectories and computed the longest OH bond length for each water molecule. Figure S6A shows these bond lengths for a representative trajectory. It is evident that only one water molecule undergoes dissociation. This behavior is consistently observed across the full ensemble of trajectories, as shown in Figure S6B, where we plot the average maximum OH bond length for each water molecule across all excited state trajectories. This approach provides a simple and reliable method for identifying the water molecule that forms the $\mathrm{HO}^{\bullet}$.
Figure S6: Identification of the hydroxyl radical. Panel A: shows the longest OH bond distances for all the water molecules in a representative trajectory on the excited state. Purple line represent the longest OH bond distance that is breaking, which can be identified within the 10 fs of the simulation. Green lines are the longest OH bonds of the remaining water. Panel B: shows the average longest OH bond distances for the molecule that is broking the bond (purple line) and for the other waters (green line). The transparent shadow represent the standard deviation. The upper axis shows the percentage of the entire ensemble of trajectories that remain on the $S_{1}$ potential at each time during the simulation.
@@
We can see from the Figure S7 exhibit a peak very close to 1, which confirm that we removed 1 electron localized in the $\mathrm{HO}^{\bullet}$ species. This result also confirm the presence of a radical rather that the ion ($\mathrm{OH}^{-}$) species in the simulation. After the remotion of the spin density around the radical we generated the modified cube file presented in the right panels in Figure S5. Using this modified cube files, we calculated the electron center and gyration radius.
@@
Using the modified cube files for the spin density (right panels in Figure S5) we can now calculate the electron center and the gyration radius using the following expressions43:
@@
where $\rho^{s}$, $r_{c}$, $\lambda_{i}$ and $r_{g}$ represent the spin density, electron center, the eigenvalues of the second moment tensor S and the gyration radius, respectively. It is important to note that the previous definitions are not strictly valid under periodic boundary conditions43. To overcome this, we follow the approach proposed in previous studies43[ref], in which the cube files are re-centered on the maximum of the spin density. This procedure ensures that the spin density decays to zero (or very small values) at the edges of the simulation box, leading in more numerically stable results.
@@
Figure S9: Radial Distribution Function of the hydrated electron. Upper panel shows the the radial distribution function $g(\textbf{r})$ between the electron and the hydrogen atoms of the waters. Lower panel shows the $g(\textbf{r})$ for the electron and the oxygen atoms of the waters. The results obtained with in the excited state with ROKS method are shown in purple. The results using the electronic ground state with Machine Learning (ML) were extracted from ref 43 are shown in green. All the results for the excited state are for the case of low gyration radius (lower than 4 Å, see main text).

### v2

@@ 4.2 Excited-State Molecular Dynamics
The excited state molecular dynamics (ESMD) simulations were initiated from the first electronic excited state ($S_{1}$), computed using the Restricted Open-shell Kohn–Sham (ROKS) method 46, 47. ROKS has proven to be a reliable approach for describing dissociative and charge-transfer states in closed-shell systems 48, 49, 50, processes that play a central role in the photochemistry of water.
@@
To propagate the system on the $S_{1}$ potential energy surface, we employed the Atomic Simulation Environment (ASE)97, a python module which is interfaced with CP2K software98. At each timestep, we calculated the energy on the electronic ground state $S_{0}$ using Density Functional Theory (DFT) to account for possible non-radiative relaxation pathways51, 49 (see below) which in our current setup is determined by examining the energy gap. The simulations were performed in the ensemble NVE using a timestep of 0.5 fs and velocity verlet algorithm to perform the nuclei propagation of the system.
@@
Both excited and ground state calculations were performed using the hybrid PBEh(40)-rVV10 functional99 in combination with the DZVP-MOLOPT-SR basis set100. This level of electronic structure theory has been successfully applied to the study of liquid water101, 102, 103, demonstrating its accuracy in reproducing ionization potential, redox potentials, electronic levels of excess electron. Carter et. al51. also reported excellent agreement with experiments for aqueous iodine solutions using the same functional within the ROKS framework. More recently the functional PBEh(40) was identified as one of the most reliable choices for describing the hydrated electron in its electronic ground state71. In our current work, the favorable agreement with reproducing the experimental UV-absorption and other quantities sensitive to the excited states of liquid water gives us confidence in the quality of our results.
@@
Each ESMD simulation was propagated until the system reached a crossing point, defined as the step where the energy gap between the first excited state ($S_{1}$) and the ground state ($S_{0}$) dropped below 0.2 eV. This criterion provides a robust and efficient way to identify non-radiative decay events and allows us to characterize the distribution of deactivation pathways, offering a lower bound on the excited-state lifetimes64, 65. A known limitation of this approach is that it does not explicitly include non-adiabatic coupling vectors (NACVs), which are key quantities for describing electronic transitions during non-radiative decay. However, a previous study on the radiolysis of water cluster in vacuum showed that NACVs increase significantly as the energy gap decreases104. This strong correlation implies that small energy gaps are reliable indicators of high non-adiabatic transition probabilities. Therefore, by using the energy gap as a decay criterion, we effectively capture the relevant non-radiative dynamics, justifying the omission of explicitly determining NACVs in our simulations.
@@
The simulations that did not exhibit a decay were continued until a maximum time of 1 ps was reached. Spin density cube files were calculated and outputted every 5 fs during all ESMD trajectories and at the step of crossing, for post-processing analysis (see section below). All the electronic calculations in the present work were performed using CP2K version 2024.298.
@@ 5.1 Coordination Numbers
Here, $i$ is the atom index, while $\kappa=10$, $\textbf{r}_{c}=1.38$ Å are parameters taken from previous works used to study the dissociation of water in the ground-state to form hydronium and hydroxide ions105, 106, 107, 108. In this manner, for an isolated hydrogen atom not attached to any water molecule, the $CN_{H}$ is zero while for a hydrogen covalently bonded to a water molecule $CN_{H}$ is 1. In the case of $CN_{O}$, it takes on a value of 3, 2 and 1, for a hydronium (ion or radical), neutral molecular water and hydroxyl species respectively.
@@ 5.2 Hydrogen Bond (HB) Network Classification
HB were identified using a geometrical criteria introduced by Kumar et. al.109: O-O distance of $\leq 3.5$ Å and H-O—O angle $\leq 30\ deg$. A water molecule with four HB is classified as part of a normal HB network, otherwise, it is labeled as a defect (see Figure 1C for categories). Sensitivity analysis to the choice of these geometric parameters is addressed in Section S3 of the Supplementary Information showing that while the hydrogen-bond criteria naturally affect the exact populations of defects, the trends and physical insights inferred are robust.
@@ 5.3 Spin-Density Analysis for Excitation Localization
where $r_{0}^{i}$ is the position of the ithi-th oxygen atom and $R=1$ Å. To estimate the degree of localization, we computed the Inverse Participation Ratio (IPR)53, 52:
@@ 5.4 HAT versus PCET Mechanism
A direct calculation of the electron center from the spin density cube files yields a position located between the $\mathrm{HO}^{\bullet}$ and the $\mathrm{H}^{\bullet}$ (Equation 4), or between the $\mathrm{HO}^{\bullet}$ and the excess electron (Equation 5). Figure S5 in the Supplementary Information illustrates this point. This result lacks a clear physical interpretation, as it does not correspond to the actual location of any species and essentially implies that the spin density is located close to the geometric center of the species involved. This problem does not arise in earlier theoretical studies28, 25, 45, 73, as those simulations typically start with a pre-inserted electron in bulk water and the formation of an additional radical species are not observed.
@@
To overcome this challenge, we first identify the $\mathrm{HO}^{\bullet}$ species by locating the water molecule with the longest OH bond in each trajectory (see Section S6 of the Supplementary Information). Once identified, we remove the contribution of the $\mathrm{HO}^{\bullet}$ to the spin density by excluding a spherical region of radius 1.5 Å centered on its oxygen atom (see Section S7 in the Supplementary Information). This step yields a modified cube file that provides a more meaningful representation of the spin density to localize the hydrated excess electron. We then compute the electron center and gyration radius from this file using the protocol described in previous studies45 (see Section S8 in the Supplementary Information). Since the cube files were saved every 5 fs and at the crossing geometry, the electron center and gyration radius for the intermediate frames were obtained using linear interpolation between consecutive cube files.
@@
Once all the species were correctly identified, we classified each trajectory according to its reaction mechanism. For this, we analyzed the final frame of each simulation (corresponding to the geometry at the point of non-radiative crossing) and we computed the distance between the electron center and the nearest hydrogen atom. If the distance was less than $0.6$ Å, the trajectory was classified as hydrogen atom transfer (HAT), otherwise it was assigned to the proton coupled electron transfer (PCET). The $0.6$ Å threshold was chosen based on the radial distribution function (RDF) between the electron and all hydrogen atoms (see Figure S9 in the Supplementary Information). Specifically, it represents the midpoint between two cases: a bound hydrogen atom (distance electron and nearest H $\sim$0 Å) and the first maximum associated with hydrogen atoms in the solvation shell of the electron (distance between electron and nearest H $\sim$1.2 Å)45, 51. A short electron–hydrogen distance indicates that both species occupy the same spatial region, consistent with the formation of a neutral hydrogen atom. In contrast, larger distances suggest that the electron remains solvated while the proton is bound to nearby water molecules, consistent with a PCET process.
@@
Author contribution
G.D.M. and A.H. conceived and designed the research. C.M. performed the ground-state MLIP simulations, and G.D.M. carried out the ESMD simulations. G.D.M., S.D.P., C.K.E., and D.D. performed the data analysis. G.D.M., C.J.M., and A.H. contributed to the interpretation and discussion of the results. G.D.M. and A.H. wrote the original manuscript with input from all authors. All authors discussed the findings and contributed to the final version of the manuscript.
Data Availability
[…]
@@
Figure S1: Initial photo-absorption in liquid water: The histogram of energy gaps calculated with ROKS is shown as light-blue bars, while the histogram of energy gaps weighted by oscillator strength from TDDFT calculations is presented as orange bars. Both theoretical spectra were scaled (while preserving relative intensities) to better visibility and comparison with the experimental absorption spectrum, depicted as a red line. The experimental spectrum was digitalized from Ref 18.
@@
The partitioning of the spin density is a key variable used in our approach to identify the water molecules that are involved in the photoexcitation in an agnostic manner. However, this approach uses an input parameter namely, the cutoff radius. Figure S2 shows the IPR histograms (same to the one presented in Figure 1B in the main text) using different parameters. We can see that the precise numbers of water molecules involved are different. Nonetheless, they all follow the same physical trends: most of the excitation are localized on a single water molecules and a small contribution of the delocalized excitation can be observed.
@@
Figure S4: Probability Distribution Function (PDF) of the time decays for PCET mechanism in purple bars and for the HAT process in light-blue bars.
@@
In this section, we address one of the key challenges discussed in the manuscript namely the proper identification of the excess electron. Figure S5 presents the spin densities for two representative frames, each corresponding to one of the two mechanisms observed in the photochemistry of water.
Figure S5: Identification of the excess electron. Original cubes files are shown in the left panels (as obtained from the CP2K software using ROKS method). The Modified cube files obtained as an output of our procedure are shown in the right panels. The center of the spin density for each kind of cube files are represented as a blue ball. We show the cube files for a selected frame exhibiting both PCET (upper panels) and HAT (lower panels) mechanisms.
As shown in the left panels of Figure S5, calculating the electron center using the original cube files results in a position (indicated by a light-blue sphere) located approximately in the middle between the $\mathrm{HO}^{\bullet}$ and the hydrated electron (top-left panel) or the $\mathrm{H}^{\bullet}$ (bottom-left panel). To address this issue, our protocol, described in the manuscript, generates a modified cube file. This improved representation, shown in the right panels of Figure S5, allows for a more physical determination of properties such as the electron center and the radius of gyration.
@@
In all trajectories, we observed that the dissociating species originates from the water molecule involved in the initial photo-absorption event. However, in cases where the excitation is delocalized (see Figure 1B in the main manuscript), it is not straightforward to determine in advance which specific molecule will dissociate. To automate this process, we first identified the two hydrogen atoms bonded to each oxygen at time $t=0$. We then calculated the longest OH bond distances for each water molecule. Figure S6A shows the longest bond distance for each water molecule for a representative trajectory. It is evident that only one water molecule undergoes dissociation. This behavior is consistently observed across the full ensemble of trajectories, as shown in Figure S6B, where we plot the average maximum OH bond length for each water molecule across all excited state trajectories. This approach provides a simple and reliable method for identifying the water molecule that forms the $\mathrm{HO}^{\bullet}$. Additional confirmation of the $\mathrm{HO}^{\bullet}$ is presented in the next section, where we quantify the amount of electron density carried by this species.
Figure S6: Identification of the hydroxyl radical. Panel A: shows the longest OH bond distances for all the water molecules in a representative trajectory on the excited state. Purple line represent the longest OH bond distance that is breaking, which can be identified within the 10 fs of the simulation. Green lines are the longest OH bonds of the remaining water. Panel B: shows the average longest OH bond distances for the molecule whose water molecule dissociates (purple line) as well as the other waters (green line). The transparent shadow represent the standard deviation. The upper axis shows the percentage of the entire ensemble of trajectories that remain on the $S_{1}$ potential at each time during the simulation.
@@
Figure S7 shows the probability density distribution function of the amount of the electron that is essentially removed from the candidate $\mathrm{HO}^{\bullet}$ species across all the ensemble of trajectories. The distribution exhibits a peak very close to 1, confirming that 1 electron is removed from a species that is the $\mathrm{HO}^{\bullet}$Ṫhis result also confirms the presence of a radical rather that the ion ($\mathrm{OH}^{-}$) in the simulation. After the removal of the spin density around the radical we generated the modified cube file presented in the right panels in Figure S5. Using this modified cube files, we determined the electron center and radius of gyration.
@@
Using the modified cube files for the spin density (right panels in Figure S5) we can now determine the electron center and radius of gyration using the following expressions45:
@@
where $\rho^{s}$, $r_{c}$, $\lambda_{i}$ and $r_{g}$ represent the spin density, electron center, the eigenvalues of the second moment tensor S and the gyration radius, respectively. It is important to note that the previous definitions are not strictly valid under periodic boundary conditions45. To overcome this, we follow the approach proposed in previous studies45, in which the cube files are re-centered on the maximum of the spin density. This procedure ensures that the spin density decays to zero (or to very small values) at the edges of the simulation box, leading to more numerically stable results.
@@
Figure S9: Radial Distribution Function of the hydrated electron. Upper panel shows the the radial distribution function $g(\textbf{r})$ between the electron and the hydrogen atoms of the waters. Lower panel shows the $g(\textbf{r})$ for the electron and the oxygen atoms of the waters. The results obtained on the excited state with the ROKS method are shown in purple. The results using a recently reported machine-learning (ML) potential from Ref45 are shown in green. All the results for the excited state are for the case of low gyration radius (lower than 4 Å, see main text).